Vector A has a magnitude of 5 units and is directed north vector b has a magnitude of 4 units and is directed west . by constructing vector diagrams find the magnitudes and direction of A+b and A-B?

Grade:12th pass

1 Answers

PASUPULETI GURU MAHESH
7 years ago
magnitude A=5
magnitude B=4
magnitude of A+B= sqrt(A2+B2+2ABcos(teeta))   (here teete=90)
magnitude of A-B= sqrt(A2+B2-2ABcos(teeta))      (here teete=90)
magnitude of A+B=magnitude of A-B
then A+B=sqrt(25+16)=root(41)
there fore magnitude of A+B=magnitude of A-B= root(41)
 
 THANK YOU